OpenStudy (anonymous):

simplify sqrt 20/sqrt 4

OpenStudy (anonymous):

any help is appreciated

OpenStudy (anonymous):

what is sqrt(4) ?

hero (hero):

\[\frac{\sqrt{4\dot\ 5}}{\sqrt{4}}\]

OpenStudy (anonymous):

2

hero (hero):

The sqrt 4's cancel

OpenStudy (anonymous):

hero summed it up already :)

hero (hero):

so the answer is just sqrt{5}

OpenStudy (anonymous):

oh okay so its just sqrt of 5 then

hero (hero):

:o)

OpenStudy (anonymous):

how do i write it though

hero (hero):

\[\sqrt{5}\]

OpenStudy (anonymous):

i meant the whole sqrt 4 being cancelled

hero (hero):

write them, then cross them out

OpenStudy (anonymous):

okay

hero (hero):